OpenStudy (anonymous):

For z = 2cis pi/6, find z^5 in rectangular form

OpenStudy (anonymous):

@wolfd.andi

OpenStudy (anonymous):

@Anon101

OpenStudy (anonymous):

Recall DeMoivre's theorem: \[\newcommand{cis}{\,\text{cis}\,}\Large \left(re^{it}\right)^n=\left(r\cis t\right)^n=r^n\cis nt=r^ne^{int}\]